No relevant resource is found in the selected language.

This site uses cookies. By continuing to browse the site you are agreeing to our use of cookies. Read our privacy policy>

Reminder

To have a better experience, please upgrade your IE browser.

upgrade

Failed to add a new NE in a ring due to LOF alarms on SL64 boards of new NE OSN7500 because switching status was 'signal failure'

Publication Date:  2012-07-25 Views:  30 Downloads:  0
Issue Description
In customers network it was required to add a new NE in existing STM-64 ring. New NE is OSN7500. All fibres of new NE are already connected with ODF. There are two methods of adding new NE in a ring, one is by manual and other is by wizard. The procedure addopted of adding new NE in a ring was by manual configuration. The details are as below:
1) Switch the traffic of NE1 and NE4 to protection as shown in attached diagram
2) Break fibre between NE1 and NE4 and add NE5 in between two NEs. Also if there are passthroughs required on NE5 then configure them first before adding it into a ring. Also create MSP in new NE and define local node ID, east node ID and west node ID according to the existing MSP ring
3) Now its time to switch the services of NE1 and NE4 to working channels
4) As switching was cleared on NE1 and NE4, instead of showing 'wait-to-restore' status on west board of NE1 and east board of NE4, it was showing 'signal-failure' status. 
Alarm Information
IN_PWR_ABN on STM64 boards of new NE
AU_LOP on STM64 boards of new NE
Handling Process
After checking optical power values of SL64 boards of new NE it was found abnormal power values on them. Reciever sensitivity of STM64 board is -17 and overload point is -1. So optical input power of STM64 board should be like: 
IN power = (overload - 5) ~ (receiver sensitivity + 3)
But input power received on both east and west STM64 of new NE were -4dBm which is high. Due to which there was signal failure. After adjusting optical power values on east and west STM64 boards upto nominal range, problem was resolved and status of ring went on 'wait-to-restore'. 
Root Cause
There are many possibilities like: 
1) Check the fibre pairs
2) Check the MSP parameters of new NE
3) Check the optical power values of STM64 boards of new NE
Suggestions
In the above case always clear AU_LOP and IN-PWR-ABN alarms first. Due to AU_LOP alarms, the status of ring goes on switching because of signal failure. 

END